### v1

# Born series for s-wave scattering length and some exact results

Abstract
In these notes the Born series for the $s$-wave scattering $a_{0}$ is calculated for a class of central potentials $V(r)$ up to sixth order in a dimensionless coupling strength $g$. Examples of exponentially decaying potentials as well truncated potentials involving a single length-scale $a$ are considered. In certain favorable cases the exact result for the $g$-dependent $s$-wave scattering length $a_{0}=A_{0}(g)a$ can be given in terms of special functions. The poles of $A_{0}(g)$ at increasing positive values of $g$ correspond to the thresholds, where $s$-wave bound-states occur successively.
1 Preparation
Consider the quantum mechanical scattering in a short-range central potential $V(r)$. The time-independent Schrödinger equation is equivalently transformed into an integral equation for the scattering wave-function [1]
$\psi_{k}(\vec{r}\,)=e^{ikz}-{m\over 2\pi\hbar^{2}}\int d^{3}r^{\prime}\,{\exp(ik|\vec{r}-\vec{r}\,^{\prime}|)\over|\vec{r}-\vec{r}\,^{\prime}|}V(r^{\prime})\psi_{k}(\vec{r}\,^{\prime})\,$, (1)
where the first part on the right hand side is the incoming planar wave in $z$-direction, and the second part is the scattered wave. Its asymptotic behavior for $|\vec{r}\,|\to\infty$ has of the well-known form [1]: complex-valued scattering amplitude $f(k,\theta)$ times an outgoing spherical wave $e^{ikr}/r$. The Born series is obtained by inserting for $\psi_{k}(\vec{r}\,^{\prime})$ first the planar wave $e^{ikz^{\prime}}$, then adding the correction obtained this way linear in $V(r^{\prime})$, and iterating the procedure. The asymptotic behavior for $|\vec{r}\,|\to\infty$ of all these contributions to the scattered wave determines the multiple scattering series [1] for the scattering amplitude $f(k,\theta)$. In the strict long-wavelength limit $k\to 0$, where the scattering amplitude $f(k,\theta)$ reduces to the $s$-wave scattering length $a_{0}=f(0,\theta)$, this series simplifies drastically, and one obtains the following expansion for $a_{0}$:
$a_{0}$ $=$ $\Big({-m\over 2\pi\hbar^{2}}\Big)\!\!\int\!\!d^{3}r_{1}V(r_{1})+\Big({-m\over 2\pi\hbar^{2}}\Big)^{2}\!\!\int\!\!d^{3}r_{1}\int\!\!d^{3}r_{2}{V(r_{1})V(r_{2})\over|\vec{r}_{1}-\vec{r}_{2}|}$ (2)
$+\Big({-m\over 2\pi\hbar^{2}}\Big)^{3}\!\!\int\!\!d^{3}r_{1}\!\!\int\!\!d^{3}r_{2}\!\!\int\!\!d^{3}r_{3}{V(r_{1})V(r_{2})V(r_{3})\over|\vec{r}_{1}-\vec{r}_{2}||\vec{r}_{2}-\vec{r}_{3}|}+\dots$
Note that the occurring integrals over a central potential $V(r_{i})$, multiplied with the ”static propagator” $|\vec{r}_{i}-\vec{r}_{j}|^{-1}$, are of the same form as electrostatic potentials produced by radially symmetric charge densities. The pertinent angular integral is in each step given by
$\int d\Omega_{i}\,{1\over|\vec{r}_{i}-\vec{r}_{j}|}={4\pi\over\text{max}(r_{i},r_{j})}\,$, (3)
[…]
In order to emphasize the expansion in a (small) parameter, it is advantageous to factor out the physical constants ($\hbar$ and $m$) and choose for the central potential $V(r)$ the ansatz:
$V(r)=-{\hbar^{2}g\over 2ma^{2}}f(r/a)\,$, (4)
with $a$ the inherent length-scale and $g$ a dimensionless coupling strength. Moreover, the profile function $f(x)$ specifies the dependence of the central potential $V(r)$ on distance $r$. In this setting the $s$-wave scattering length is necessarily a multiple of the length-scale $a$, namely $a_{0}=A_{0}a$. The expansion of the dimensionless quantity $A_{0}$ in powers of the dimensionless coupling strength $g$ reads
$A_{0}(g)=c_{1}g+c_{2}g^{2}+c_{3}g^{3}+c_{4}g^{4}+c_{5}g^{5}+c_{6}g^{6}+\dots\,$, (5)
where the first few coefficients are computed as
$c_{1}=\int_{0}^{\infty}\!\!dx\,x^{2}f(x)\,,\quad c_{2}=\int_{0}^{\infty}\!\!dx\,x^{2}f(x)\Phi(x)\,,\quad c_{3}=\int_{0}^{\infty}\!\!dx\,x^{2}f(x)[\Phi(x)]^{2}\,$, (6)
$c_{4}=\int_{0}^{\infty}\!\!\!\!dx_{1}\!\int_{0}^{\infty}\!\!\!\!dx_{2}\,{x_{1}^{2}x_{2}^{2}\over\text{max}(x_{1},x_{2})}\Phi(x_{1})f(x_{1})f(x_{2})\Phi(x_{2})=2\!\int_{0}^{\infty}\!\!\!\!dx_{1}\!\int_{0}^{x_{1}}\!\!\!\!dx_{2}\,x_{1}x_{2}^{2}\Phi(x_{1})f(x_{1})f(x_{2})\Phi(x_{2}).$
Here, $\Phi(x)$ is the ”electrostatic potential” produced by the ”charge density” $f(x)$:
$\Phi(x)=\int_{0}^{\infty}\!dx^{\prime}{x^{\prime 2}\over\text{max}(x,x^{\prime})}f(x^{\prime})=\int_{0}^{x}\!dx^{\prime}\,{x^{\prime 2}\over x}f(x^{\prime})+\int_{x}^{\infty}\!dx^{\prime}\,x^{\prime}f(x^{\prime})\,.$ (7)
Higher coefficients $c_{j}$ involve further factors of $x_{i}^{2}f(x_{i})/\text{max}(x_{i},x_{j})$ in a multiple integral. Note that the increasing powers of the prefactor in the Born series ($m/2\pi\hbar^{2}$) and in the potential ($\hbar^{2}/2m$), together with $4\pi$ from the angular integrations, have compensated each other at every order.
2 Exponentially decaying potentials
In this section the Born series $A_{0}=\sum_{j=1}^{6}c_{j}g^{j}+\dots$ is computed for a variety of central potentials $V(r)$ with an exponential decay. The familiar Yukawa potential is physically well motivated in terms of an underlying one-boson exchange mechanism. The pertinent profile function $f(x)$ and ”electrostatic potential” $\Phi(x)$ are in this case
$f(x)={e^{-x}\over x}\,,\qquad\Phi(x)={1-e^{-x}\over x}\,.$ (8)
With this input one derives for the Born series up to sixth order:
$A_{0}(g)=g+{g^{2}\over 2}+g^{3}\ln{4\over 3}+g^{4}\ln{32\over 27}+0.100941276\,g^{5}+0.06006048\,g^{6}+\dots$ (9)
Note that in a Feynman diagrammatic picture, where the Yukawa potential stems from a massive one-boson exchange, order $g$ corresponds to tree-level, order $g^{2}$ to one-loop, order $g^{3}$ to two-loops, order $g^{4}$ to three-loops, and so on. The two-loop coefficient $c_{3}=\ln(4/3)=0.28768207$ has also been found in a momentum-space calculation of twice-iterated one-boson exchange in ref.[2], while the three-loop coefficient $c_{4}=\ln(32/27)=0.16989904$ seems to be a novel result. It is quite astonishing that even the four-loop coefficient $c_{5}=0.1009412761$ can be given in exact form as
$c_{5}={\pi^{2}\over 6}+(10+3\ln 2-2\ln 3)\ln 2-{9\over 2}\ln 3-{5\over 4}\ln 5+2\text{Li}_{2}\Big(\!-{1\over 2}\Big)-2\text{Li}_{2}\Big(\!-{1\over 3}\Big)+2\text{Li}_{2}\Big(\!-{2\over 3}\Big)\,$, (10)
[…]
The series for $A_{0}(g)$, with its first six terms written in eq.(8), has a finite radius of convergence $(|g|<g_{1})$, that is determined as that strength parameter $g_{1}$, where the threshold for the first $s$-wave bound-state in the Yukawa potential is reached. This particular value $g_{1}$ can be computed by solving numerically the second-order differential equation for the radial wave-function (multiplied with $x$):
$u^{\prime\prime}(x)+g{e^{-x}\over x}u(x)=0\,$, (11)
subject to a linear behavior $u(x)\sim x$ at small $x$. Note that $u^{\prime\prime}(x)+gx^{-1}u(x)=0$ has the regular solution in the form of a Bessel-function $u(x)=\sqrt{g\,x}J_{1}(2\sqrt{g\,x})=g\,x+{\cal O}(x^{2})$, which also displays a linear behavior. Once the critical strength $g_{1}$ is reached, the asymptotic behavior of $u(x)$ for large $x$ turn into a constant. The convergence of this process is monitored by inspecting, how the first derivative $u^{\prime}(x)$ approaches zero for large $x$, when fine-tuning more and more digits of the strength parameter $g$. As a result of this numerical exercise, one finds that the first threshold $g_{1}$ for an $s$-wave bound-state in the Yukawa potential lies at the value:
$g_{1}=1.7443138\,.$ (12)
Next, we consider a purely exponentially decreasing potential $V(r)\sim-g\,e^{-r/a}$ with the profile function
$f(x)=e^{-x}\,,\qquad\Phi(x)={2\over x}-\Big(1+{2\over x}\Big)e^{-x}\,$, (13)
and on this basis, one derives the following Born series up to sixth order
$A_{0}=2\,g+{5\over 4}\,g^{2}+{23\over 27}\,g^{3}+{677\over 1152}\,g^{4}+{7313\over 18000}\,g^{5}+{218491\over 777600}\,g^{6}+\dots$ (14)
For the exponential potential $\sim-g\,e^{-x}$ a little miracle happens, since (at $k=0$) the second-order linear homogeneous differential equation for the radial wave-function (multiplied with $x$):
$u^{\prime\prime}(x)+g\,e^{-x}u(x)=0\,$, (15)
possesses analytical solutions in the form of a Bessel-function $u_{1}(x)=J_{0}(2\sqrt{g}\,e^{-x/2})$ and a Neumann-function $u_{2}(x)=N_{0}(2\sqrt{g}\,e^{-x/2})$ [3] of index 0. The proper linear combination is of these two is found by demanding $u(0)=0$ and an asymptotic behavior $u(x)\simeq x+A_{0}$ for $x\to\infty$. The latter asymptotic behavior $u(x)\simeq x(1+A_{0}/x)$ is provided by the logarithmic singularity of the Neumann-function $N_{0}(z)=2[\gamma_{E}+\ln(z/2)]/\pi+{\cal O}(z^{2}\ln z)$ near $z=0$. Putting the pieces together, the resulting exact expression for the $s$-wave scattering length in the exponential potential reads:
$A_{0}(g)={\pi N_{0}(2\sqrt{g})\over J_{0}(2\sqrt{g})}-\ln g-2\gamma_{E}\,,\quad\text{for}\,\,\,g>0\,$,
$A_{0}(g)=-{2K_{0}(2\sqrt{-g})\over I_{0}(2\sqrt{-g})}-\ln(-g)-2\gamma_{E}\,,\quad\text{for}\,\,\,g<0\,$, (16)
Figure 1: Scattering length $A_{0}=a_{0}/a$ versus $g$ for the exponential potential $V(r)\sim-g\,e^{-r/a}$.
with $\gamma_{E}=0.577215665$ the Euler-Mascheroni number. In the second formula (for $g<0$), $K_{0}$ and $I_{0}$ are modified Bessel-functions [3] that arise from the continuation of $N_{0}$ and $J_{0}$ to imaginary argument. The behavior of $A_{0}(g)$ is shown in Fig. 1 for both positive and negative values of the strength parameter 20<g<20-20<g<20. One observes poles at positive values of $g$, which mark the thresholds for successive $s$-wave bound-states in the attractive exponential potential. According to exact formula in eq.(16), these thresholds can be identified with the zeros $z_{j}$ of the Bessel-function $J_{0}(z)$ through the relation $g_{j}=z_{j}^{2}/4$. The first six threshold values read
$g_{1}\!=\!1.445796\,,\quad g_{2}\!=\!7.617816\,,\quad g_{3}\!=\!18.72175\,,\quad g_{4}\!=\!34.76007\,,\quad g_{5}\!=\!55.73308\,,\quad g_{6}\!=\!81.64084\,.$ (17)
Note that the numerical method, employed before to construct $g_{1}$ for the Yukawa potential, reproduces $g_{1}=1.4457965$ for the profile $f(x)=e^{-x}$.
As a further example, we choose an exponentially decreasing potential with a soft core:
$f(x)=x\,e^{-x}\,,\qquad\Phi(x)={6\over x}-\Big({6\over x}+4+x\Big)e^{-x}\,$, (18)
and derive the Born series up to sixth order
$A_{0}(g)=6\,g+{33\over 4}\,g^{2}+{5873\over 486}\,g^{3}+17.835277g^{4}+26.351768\,g^{5}+38.941427\,g^{6}+\dots\,$, (19)
where all expansion coefficients are in fact rational numbers. The radius of convergence $|g|<g_{1}$ of the series for $A_{0}(g)$ written in eq.(19) is determined by searching the threshold for the first $s$-wave bound-state in the potential $V(r)\sim-g\,x\,e^{-x}$. One finds with high numerical accuracy the value
$g_{1}=0.676681997\,.$ (20)
Next, we choose an exponentially decreasing potential with a very soft core, taking the profile function
$f(x)=x^{2}e^{-x}\,,\qquad\Phi(x)={24\over x}-\Big({24\over x}+18+6x+x^{2}\Big)e^{-x}\,$, (21)
and derive the following Born series up to sixth order
$A_{0}(g)=24\,g+{837\over 8}\,g^{2}+479.53818\,g^{3}+2209.6103\,g^{4}+10188.272\,g^{5}+46981.234\,g^{6}+\dots\,$, (22)
where all expansion coefficients are (large) rational numbers. The radius of convergence $|g|<g_{1}$ of the series for $A_{0}(g)$ in eq.(22) is again determined by searching the threshold for the first $s$-wave bound-state in the potential $V(r)\sim-g\,x^{2}e^{-x}$. One finds with high numerical accuracy the value
$g_{1}=0.216855258\,.$ (23)
Finally, one can consider an exponentially decreasing potential with a more singular behavior at small distances:
$f(x)={e^{-x}\over x^{2}}\,,\qquad\Phi(x)={1-e^{-x}\over x}+\Gamma(0,x)\,$, (24)
where $\Gamma(0,x)=\int_{x}^{\infty}\!d\xi\,\xi^{-1}e^{-\xi}=-\ln x-\gamma_{E}+x-x^{2}/4+{\cal O}(x^{3})$ denotes the incomplete $\Gamma$-function [3]. With this input one derives the following Born series up to sixth order:
$A_{0}(g)=g+g^{2}\ln 4+2.98036548\,g^{3}+7.68193343\,g^{4}+21.86973\,g^{5}+66.2866\,g^{6}+\dots$ (25)
The profile $f(x)=e^{-x}/x^{2}$ is too singular near the origin and therefore the coupling strength $g$ must be restricted to negative and small positive values $g<g_{\text{crit}}=1/4$. This feature concerning an instability is further elaborated on in the next section, when analyzing the truncated $1/r^{2}$-potential.
3 Truncated potentials
A well-known textbook problem [1] is the quantum mechanical scattering off a spherical potential well, represented by the profile
$f(x)=\theta(1-x)\,,\qquad\Phi(x)={1\over 2}-{x^{2}\over 6}\,$, (26)
and now the knowledge of the ”electrostatic potential” $\Phi(x)$ is needed only in the inside region $0<x<1$. With little effort, one obtains the following Born series for the $s$-wave scattering length in a spherical potential well:
$A_{0}(g)={g\over 3}+{2\over 15}\,g^{2}+{17\over 315}\,g^{3}+{62\over 2835}g^{4}+{1382\over 155925}\,g^{5}+{21844\over 6081075}\,g^{6}+\dots$ (27)
The analytical solution to the differential equation for the radial wave-function $u^{\prime\prime}(x)+g\,u(x)=0$ (in the region $0<x<1$) subject to the boundary condition $u(0)=0$ is either a sine-function, $\sin(\sqrt{g}x)$, or a sine-hyperbolic function, $\sinh(\sqrt{-g}x)$. After matching the logarithmic derivative of $u(x)$ at $x=1$ to that of the linear form $u(x)=x+A_{0}$ outside, one obtains the generic formula $A_{0}=u(1)/u^{\prime}(1)-1$ for the $s$-wave scattering length in any truncated potential. In the present case one reproduces the well-known result
$A_{0}(g)=-1+{1\over\sqrt{g}}\tan\sqrt{g}\,,\quad\text{for}\,\,\,g>0\,$,
$A_{0}(g)=-1+{1\over\sqrt{-g}}\tanh\sqrt{-g}\,,\quad\text{for}\,\,\,g<0\,$, (28)
and, as a nice check, the Taylor expansion of the tangent-function confirms the series in eq.(27). The analytical result for $A_{0}(g)$ written in eq.(28) is shown in Fig. 2 in the region 40<g<62-40<g<62. The poles of $A_{0}(g)$ at the values $g_{j}=\pi^{2}(j-1/2)^{2}$ (where $\tan\sqrt{g_{j}}$ diverges) correspond to the well-known thresholds for the occurrence of successive $s$-wave bound-states in a spherical potential well. The first six values of this infinite sequence read
$g_{1}\!=\!2.467401\,,\quad g_{2}\!=\!22.20661\,,\quad g_{3}\!=\!61.68503\,,\quad g_{4}\!=\!120.9027\,,\quad g_{5}\!=\!199.8595\,,\quad g_{6}\!=\!298.5555\,.$ (29)
Figure 2: $A_{0}$ as a function of $g$ for the spherical well potential.
It is also interesting to study the truncated $1/r^{2}$-potential with the profile function
$f(x)=\Big({1\over x^{2}}-1\Big)\theta(1-x)\,,\qquad\Phi(x)={1\over 2}+{x^{2}\over 6}-\ln x\,$, (30)
and this input leads to the following Born series for the $s$-wave scattering length up to sixth order
$A_{0}(g)={2\over 3}\,g+{56\over 45}\,g^{2}+{14312\over 4725}\,g^{3}+8.364150\,g^{4}+24.881527\,g^{5}+77.758378\,g^{6}+\dots$ (31)
The differential equation for the radial wave-function in the inner region $(0<x<1)$:
$u^{\prime\prime}(x)+g\Big({1\over x^{2}}-1\Big)u(x)=0\,$, (32)
possesses a regular solution in the form of a Bessel-function $u(x)=\sqrt{x}J_{\nu}(\sqrt{-g}x)$ with index $\nu=\sqrt{1/4-g}$. Note that the strength parameter $g$ has to be constrained to $g\leq 1/4$ in order to evade the instability of the quantum mechanical problem with a too attractive $1/r^{2}$-potential. For comparison the regular solution of $u^{\prime\prime}(x)+g\,x^{-2}u(x)=0$ is for all $g\leq 1/4$ given by $u(x)=x^{\nu+1/2}$. Under this restriction on $g$ the logarithmic derivative $u^{\prime}(x)/u(x)$ at $x=1$ determines the dimensionless $s$-wave scattering length as $A_{0}=u(1)/u^{\prime}(1)-1$. In the present case this leads to the expression:
$A_{0}(g)=\bigg\{{1+\sqrt{1-4g}\over 2}-\sqrt{-g}\,{J_{\nu+1}(\sqrt{-g})\over J_{\nu}(\sqrt{-g})}\bigg\}^{-1}-1\,,\qquad\nu={\sqrt{1-4g}\over 2}\,$, (33)
where second term in curly brackets, $\sqrt{-g}J_{\nu+1}(\sqrt{-g})/J_{\nu}(\sqrt{-g})$, means $\sqrt{g}I_{\nu+1}(\sqrt{g})/I_{\nu}(\sqrt{g})$ for $0<g<1/4$. The exact result for $A_{0}(g)$ written in eq.(33) is shown in Fig. 3 for the parameter range 5<g<1/4-5<g<1/4. As already visible in Fig. 2, and valid for any truncated potential, the scattering length approaches with increasing repulsion the hard-core value: $\lim_{g\to-\infty}A_{0}(g)=-1$. On the positive side the boundary value, before instability sets in, is given by
$A_{0}(1/4)={I_{0}(1/2)-I_{1}(1/2)\over I_{0}(1/2)+I_{1}(1/2)}=0.6095845\,.$ (34)
Next, we consider the truncated Coulomb-potential with profile function
$f(x)=\Big({1\over x}-1\Big)\theta(1-x)\,,\qquad\Phi(x)={1-x\over 2}+{x^{2}\over 6}\,$, (35)
Figure 3: $A_{0}$ as a function of $g$ for the truncated $1/r^{2}$-potential, restricted to the range $g\leq 1/4$.
Figure 4: $A_{0}$ as a function of $g$ for the truncated Coulomb-potential.
and derive the following Born series up to sixth order
$A_{0}(g)={g\over 6}+{g^{2}\over 20}+{241\over 15120}\,g^{3}+{233\over 45360}\,g^{4}+{g^{5}\over 602.7393}+{g^{6}\over 1865.2312}+\dots$ (36)
Now, the differential equation $u^{\prime\prime}(x)+g(x^{-1}-1)u(x)=0$ for the radial wave-function possesses a regular solution of the form $u(x)=x\,e^{-\sqrt{g}\,x}\,_{1}F_{1}(1-\sqrt{g}/2,2;2\sqrt{g}\,x)$, including a confluent hypergeometric series [3]. Based on this solution the exact expression for the $s$-wave scattering length in the truncated Coulomb-potential $A_{0}(g)=u(1)/u^{\prime}(1)-1$ takes the form:
$A_{0}(g)=\bigg\{1-\sqrt{g}+{(\sqrt{g}-g/2)\,\,_{1}F_{1}(2-\sqrt{g}/2,3;2\sqrt{g})\over{}_{1}F_{1}(1-\sqrt{g}/2,2;2\sqrt{g})}\bigg\}^{-1}-1\,.$ (37)
By solving the equation, that the numerator of the $g$-dependent function in curly brackets is zero, one determines the threshold for successive $s$-wave bound-states in the truncated Coulomb-potential. The first six zeros are computed as:
$g_{1}\!=\!3.094169\,,\quad g_{2}\!=\!13.88101\,,\quad g_{3}\!=\!32.62548\,,\quad g_{4}\!=\!59.35213\,,\quad g_{5}\!=\!94.068586\,,\quad g_{6}\!=\!136.77831\,.$ (38)
The expression for $A_{0}(g)$ written in eq.(37) is plotted in Fig. 4 in the range 33<g<33-33<g<33.
Figure 5: $A_{0}$ as a function of $g$ for the truncated oscillator potential.
Figure 6: $A_{0}$ as a function of $g$ for the truncated linear potential.
A further interesting example is the truncated oscillator potential with the profile function
$f(x)=(1-x^{2})\theta(1-x)\,,\qquad\Phi(x)={1\over 4}-{x^{2}\over 6}+{x^{4}\over 20}\,$, (39)
and one derives the following Born series up to sixth order
$A_{0}(g)={2\over 15}\,g+{8\over 315}\,g^{2}+{1432\over 289575}\,g^{3}+{g^{4}\over 1036.0372}+{g^{5}\over 5306.4658}+{g^{6}\over 27178.112}+\dots$ (40)
The solution of the differential equation $u^{\prime\prime}(x)+g(1-x^{2})u(x)=0$ can be found in terms of parabolic cylinder functions $D_{\nu}(\sqrt{\pm 2}g^{1/4}x)$ [3] with respective indices $\nu=(-1\pm\sqrt{g})/2$. From the pertinent linear combination which satisfies $u(0)=0$, one gets through the formula $A_{0}(g)=u(1)/u^{\prime}(1)-1$ a very lengthy expression that will not be presented here. The pertinent numerical result is shown in Fig. 5 for the parameter range 110<g<110-110<g<110. The first six poles of $A_{0}(g)$ corresponding to thresholds for $s$-wave bound-states in the truncated oscillator potential lie at:
$g_{1}\!=\!5.121669\,,\quad g_{2}\!=\!39.66084\,,\quad g_{3}\!=\!106.2492\,,\quad g_{4}\!=\!204.8561\,,\quad g_{5}\!=\!335.4732\,,\quad g_{6}\!=\!498.0971\,.$ (41)
As a last example we consider the truncated linear potential with the profile function
$f(x)=(1-x)\theta(1-x)\,,\qquad\Phi(x)={1-x^{2}\over 6}+{x^{3}\over 12}\,$, (42)
and derive the following Born series up to sixth order:
$A_{0}(g)={g\over 12}+{13\over 1260}\,g^{2}+{17\over 12960}\,g^{3}+{g^{4}\over 5977.6798}+{g^{5}\over 46852.232}+{g^{6}\over 367200.57}+\dots$ (43)
The solution of the differential equation $u^{\prime\prime}(x)+g(1-x)u(x)=0$ for the radial wave-function (in the region $0<x<1$) subject to the boundary condition $u(0)=0$ is given by a linear combination of Airy-functions [3]. Its logarithmic derivative at $x=1$ leads to the following exact expression for the $s$-wave scattering length in a truncated linear potential:
$A_{0}(g)={\big[\sqrt{3}\text{Ai}(-g^{1/3})-\text{Bi}(-g^{1/3})\big]\Gamma(1/3)\over(3g)^{1/3}\big[\sqrt{3}\text{Ai}(-g^{1/3})+\text{Bi}(-g^{1/3})\big]\Gamma(2/3)}-1\,$, (44)
which allows to confirm the Taylor series expansion written in eq.(43). The numerical result for $A_{0}(g)$ written in eq.(44) is shown in Fig. 6 over the large parameter range 150<g<150-150<g<150. The poles of $A_{0}(g)$ at positive $g$ demarcate the thresholds for successive $s$-wave bound-states in the truncated linear potential. The first six zeros of the pertinent equation, $\sqrt{3}\text{Ai}(-g^{1/3})+\text{Bi}(-g^{1/3})=0$, are computed as
$g_{1}\!=\!7.837347\,,\quad g_{2}\!=\!55.97703\,,\quad g_{3}\!=\!148.5083\,,\quad g_{4}\!=\!285.4502\,,\quad g_{5}\!=\!466.8047\,,\quad g_{6}\!=\!692.5722\,.$ (45)
4 Further solvable potential scattering problem
In passing we present and discuss another solvable scattering problem, where the potential has the shape of a squared Lorentz-curve, i.e. the profile function is
$f(x)={1\over(1+x^{2})^{2}}\,,\qquad\Phi(x)={\arctan x\over 2x}\,.$ (46)
The pertinent second-order differential equation for the radial wave-function
$u^{\prime\prime}(x)+{g\over(1+x^{2})^{2}}u(x)=0\,$, (47)
possesses the analytical solution $u(x)=\sqrt{1+x^{2}}\sin(\sqrt{1+g}\arctan x)$, subject to the boundary condition $u(0)=0$. From its asymptotic behavior $u(x)\simeq const(x+A_{0})$ for $x\to\infty$, one can deduce with little effort the following exact result for the dimensionless $s$-wave scattering length:
$A_{0}(g)=-\sqrt{1+g}\cot\Big({\pi\over 2}\sqrt{1+g}\Big)\,,\quad\text{for}\,\,\,g>-1\,$,
$A_{0}(g)=-\sqrt{-1-g}\coth\Big({\pi\over 2}\sqrt{-1-g}\Big)\,,\quad\text{for}\,\,\,g<-1\,$, (48)
with the intermediate value $A_{0}(-1)=-2/\pi=-0.63661977$. The behavior of $A_{0}(g)$ written in eq.(48) is shown in Fig. 7 over the wide parameter range 100<g<100-100<g<100. The poles of $A_{0}(g)$, which demarcate the thresholds for $s$-wave bound-states in a potential of squared Lorentz shape, lie at coupling strengths $g_{j}=4j^{2}-1$. This sequence of integer values reads up to $j=6$:
$g_{1}=3\,,\qquad g_{2}=15\,,\qquad g_{3}=35\,,\qquad g_{4}=63\,,\qquad g_{5}=99\,,\qquad g_{6}=143\,.$ (49)
The Taylor expansion of $A_{0}(g)$ written in eq.(48) gives the following Born series up to sixth order
$A_{0}(g)$ $\!\!=\!\!$ $g\,{\pi\over 4}+g^{2}\,{\pi\over 16}+g^{3}\,{\pi(\pi^{2}-6)\over 192}+g^{4}\,{\pi(15-\pi^{2})\over 768}+g^{5}\,{\pi(2\pi^{4}+5\pi^{2}-210)\over 15360}+g^{6}\,{\pi(105-\pi^{4})\over 10240}\dots$
$\!\!=\!\!$ $0.7853982\,g+0.1963495\,g^{2}+{g^{3}\over 15.793733}+{g^{4}\over 47.649735}+{g^{5}\over 143.10164}+{g^{6}\over 429.39433}+\dots$
It is a good check to reproduce the occurring expansion coefficients $c_{1},\dots,c_{6}$ by evaluating the multiple integrals in eq.(6) with the profile functions $f(x)$ and $\Phi(x)$ written in eq.(46).
Figure 7: $A_{0}$ as a function of $g$ for the potential of squared Lorentz shape.
This concludes our presentation of results for the $s$-wave scattering length $A_{0}(g)$ for selected classes of short-range central potentials $V(r)\sim-g\,f(x)$. Of course, one could continue these analytical and numerical calculations for countless other profiles functions $f(x)$. A closed form expression for $A_{0}(g)$ can be expected to exist only in rare cases.